OpenStudy (anonymous):

write the expression as a single logarithm 1. 2 logbq + 8 logbt

OpenStudy (ranga):

Is b the base or are we taking the log of the product of b and q?

OpenStudy (anonymous):

\[2\log _{b}q +8 \log _{b}q\] did this help im not sure what it stands for

OpenStudy (anonymous):

\[\log _{b}(q ^{2}t ^{8})\] this is what i think it is

OpenStudy (ranga):

Okay, b is the base. \[\Large 2\log _{b}q +8 \log _{b}q = \log _{b}q^2 + \log _{b}q^8 = \log _{b}(q^2 * q^8) = \log _{b}(q^{10})\]

OpenStudy (ranga):

The last part is\[\Large \log_{b} (q^{10})\]

OpenStudy (anonymous):

thank you!

OpenStudy (ranga):

This just makes use of a couple of logarithmic identities: n * log(A) = log(A)^n and log(A) + log(B) = log(AB) You are welcome.